hi Algebra.Com tutors,
thanks for your mathematical services. I have been dealing with permutations for a while, so I'm a genius at solving them, I believe.
however, the problem challenging me is working them practically! I solved 3 P 2 (=6) by drawing a table. But 5 P 3 (=60) and 10 P 3 (=720) failed completely!
Can you help me with the true table proving the results above?
I look forward to hear from you.


A. S. B. H.

Answer by rothauserc(4718)

You are dealing with permutations where you have n itmes
:
For example 3 P 2 means you have three items from which you are making selections
:
3 P 2 = 3 * (3-1) = 6 (for the first choice you have 3 to pick from, for second choice you have 2 to pick from)
:
5 P 3 = 5 * (5-1) * (5-2) = 60
:
10 P 3 = 10 * (10-1) * (10-2) = 720